1. What does the expression s/4 represent?
2. What is the solution to the equation s/4=14?
3. Consider the pan balance. What would happen if you replaced the 8-ounce weight with a 10-ounce weight?
4. Supposed you change one of the 4-ounce weights to a 7-ounce weight. How much weight would you have to add on the right side to make the pans balance? How do I explain it?
Answer #1

Hello,

Sol 1) Here, s is any variable. So s can take any value. The expression s/4 represents that variable s is divided by 4 and that is the answer.

Sol 2) s/4 = 14

Here, The solution to this equation is s = 14*4 = 56.

Hence, s = 56 is the solution of the equation.

Sol 3) In a pan balance, you need to put weights on both sides. Hence, it is level when there is 8 ounce weight on both sides. But when we replace one 8 ounce weight with 10 ounce weight, then it weighs more on 10 ounce weight and it gets down on its side.

Sol 4) If i replaced 4 ounce weight with 7 ounce weight on one side, then i will need to add 3 ounce weight on the other side to make the pans balance. This is because 3 ounce got increased on the side where we replace 4 with 7 ounce weight. Now to balance it, we need to add 3 ounce weight on other side.

Thanks